It took a crew 4h 30m to row 8km upstream and back again. If the rate of flow of the stream was 7 km/h, what was the rowing speed of the crew in still water?
Answer by jorel555(1290)

If the distance was 8 km each way, then:
8/(x+7)+8/(x-7)=9/2 hrs.
16/(x+7)+16/(x-7)=9
16(x-7)+16(x+7)=9(x^2-49)
16x-112+16x+112=9x^2-441
9x^2-32x-441=0
Factoring 9x^2-32x-441 we get (9x+49)(x-9). So our answer is 9 km/h as the speed of the rowing crew in still water!!!!!!!